RunPlanDaily

From PREX Wiki
Jump to: navigation, search



PREX Main << Run Coordinator Tools

Global Run Plan

Current Daily Run Plans

The current day is indicated by a color.

Day, Date Shift Tasks (Responsible Party) Status
Monday, August 12 Owl Production running at 70uA with
Day
  1. Production running at 70uA
  2. Commissioning targets required for Transverse measurements include raster size calibrations
Swing Production running at 70uA
Tuesday, August 13 Owl Production running at 70uA
Day
  1. Planned accelerator downtime:Beam studies
  2. Preparations for Transverse Polarized Measurement
Swing Transverse Polarized Measurement
Wednesday, August 14 Owl Transverse Polarized Measurement
Day
  1. Transverse Polarized Measurement
  2. Wein Flip back to `Right'
  1. Transverse work done at ~10:30.
  2. Hand things off to OPS for beam studies and Wien flip.
Swing
  1. Establish PQB after Wien flip
  2. PITA scan
Production running at 70uA
  1. Caryn checks PQB (at ILD). Done by ~16:30
  2. PITA scan ~18:45
  3. Moved to D5-208Pb8-D6 target ~20:45
Thursday, August 15 Owl Production running at 70uA
Day Production running at 70uA
Swing Production running at 70uA
  1. Target out for Compton background studies
  2. Production back on Pb#8
Friday, August 16 Owl Production running at 70uA Beam off at 7 AM to cool down the hall prior to Compton work
Day
  1. Work in the hall for to replace Compton fiber amp
  2. Jessie will work on Tungsten Collimator flow reedback issue
  1. Jack added additional shielding around the hall
  2. Dave, Amali and Caryn worked on the laser
  3. CHL problems starting at 12:03 means no opportunity of beam during shift
Swing Production running at 70uA No beam due to CHL and Hall B restore
Saturday, August 17 Owl Production running at 70uA No beam due to CHL and Hall B restore
Day Production running at 70uA
  1. Beam returns around 8:30 AM
Swing Production running at 70uA
Sunday, August 18 Owl Production running at 70uA
Day Moller measurement
Swing Production running at 70uA
  1. Moller measurement concluded
  2. 70 uA beam restored to hall
  3. Production running
Monday, August 19 Owl Production running at 70uA
Day Production running at 70uA
Swing Production running at 70uA
Tuesday, August 20 Owl Production running at 70uA
Day
Swing

Previous Daily Run Plans

Instructions for RC

  • Let the day shift know what to do while you attend the MCC meeting
  • Copy the daily run plans for the previous week to the Run plan history
  • Copy the Run Plan Template table to this page and change the dates
  • Update the plan above (if necessary) after the MCC meeting